OpenStudy (jjuden):

If m varies directly with n, and m = 6 when n = 9, what is the value of n when m = 8?

OpenStudy (jjuden):

a) n = 24 b) n = 12 c) n = 16/3 d) n = 27/4

OpenStudy (anonymous):

M = 6 when N = 9 Half of 6 is 3. And 6 + 3 = 9. So can you do the same thing when M = 8? @jjuden

OpenStudy (anonymous):

\[\huge m=kn (k = non-zero constant)\] \[\huge 6=9k\] \[\huge k=\frac{ 6 }{ 9 }\] Same thing you can do it for the other part

OpenStudy (anonymous):

Half of 8 is 4. And 8 + 4 = 12. You could've also done: 9 / 6 = 1.5 And 8 x 1.5 = 12.

OpenStudy (anonymous):

@jjuden

OpenStudy (anonymous):

\[n = 12\]

OpenStudy (anonymous):

Because \[m =k \times n\] \[6 = k \times 9\] \[k = \frac{ 6 }{ 9 }=\frac{ 2 }{ 3 }\] So \[8 = \frac{ 2 }{ 3 }\times n\] \[n = \frac{ 8\times 3 }{ 2 } = \frac{ 24 }{ 2 }= 12\]